# Language Models Do Not Have Human-Like Working Memory

### v2

@@ 5 Math Magic
Let’s play a game before presenting the final part of our experiments. Think of a three-digit number in which the hundreds and units digits differ (e.g., abc). Reverse the digits to form a new number (e.g., abc $\rightarrow$ cba), and subtract the smaller of the two numbers from the larger. Then, reverse the result—if it is a two-digit number, prepend a zero (e.g., 67 $\rightarrow$ 067 $\rightarrow$ 760). Add this reversed number to the previous result What if we can tell that your result is 1089? It is math who tells us the secret. Such magic, often referred to as “math magic” or “math mentalism,” rely on carefully constructed transformations that yield predictable outcomes, allowing magicians to perform mind-reading.
To perform this task, humans must internally generate or retain numerical information (e.g., remember a poker card) and manipulate it mentally without external cues. This process relies on working memory, a cognitive function that LLMs may lack. Without this capacity, LLMs are prone to errors in mathematical tricks, failing to produce the intended outcomes. Accordingly, we can design complex math magics to evaluate whether LLMs possess working memory–like capabilities.
@@
Since the 1089 game is well known and likely included in the training data of LLMs, our preliminary experiments show that LLMs can recognize the game and thus invalidate the evaluation. Therefore, to more effectively assess LLMs’ multi-step mental manipulation, we select a more complicated math magicbased on the Josephus Problem Schumer (2002). In this task, participants are asked to imagine four numbers and perform a sequence of operations, including duplication, rotation, and removal. The procedure is shown in Table 6. Ultimately, only two numbers remain, and mathematical constraints guarantee they are identical. In our experiment, LLMs are prompted to secretly think of four numbers and mentally execute the sequence of randomized operations. We measure the proportion of trials (out of 150) in which the LLM completes the task correctly.
@@
Findings.
Table 7a reports the accuracy of prompting models to output the two numbers directly. Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family. This finding aligns with results from the number guessing game shown in Table 1, where LLaMA models generate more realistic distributions than other models. Together, these consistent results suggest that current LLMs generally lack the working memory required for such tasks.
@@
We further examine whether CoT prompting improves performance on this task. Table 7b presents the results of prompting models to reason step by step, as well as the performance of o1-like long reasoning models (LRMs). Base models prompted to reason step-by-step achieve 10–30% accuracy—substantially higher than without CoT. DeepSeek-R1 attains 100% accuracy, and other LRMs also perform well, indicating that extended token-level reasoning substantially improves accuracy. However, consistent with our number-guessing experiment, models exhibit a strong preference for the number seven, which appears to influence performance—especially within the o-series. Among o1-Mini’s 75 correct predictions, 50 (66.7%) select 7; for o3-Mini and o4-Mini, the counts are 68/145 (46.9%) and 37/54 (68.5%), respectively. These findings suggest that models may not fully solve the task but instead exploit biased number preferences: notably, o3-Mini—being least likely to guess 7—outperforms the other two o-series models. In conclusion, current LLMs struggle with a class of tasks demanding internal object imagination and mental manipulation, relying on an effective working memory.
6 Conclusion
@@
In this study, we present three carefully designed experiments to investigate whether LLMs possess human-like working memory for processing information and generating responses. Across all experiments, the results reveal a consistent pattern: LLMs do not exhibit behavior indicative of a functional working memory. They fail to internally represent and manipulate transient information across multiple reasoning steps, relying instead on the immediate prompt context. Even advanced prompting strategies, such as CoT prompting, yield only marginal improvements on tasks requiring internal state management. This limitation leads to unrealistic responses, self-contradictions, and an inability to perform mental manipulations. We hope these findings encourage future research on integrating working memory mechanisms into LLMs to enhance their reasoning capabilities and human-likeness.
The Use of Large Language Models
LLMs were employed in a limited capacity for writing optimization. Specifically, the authors provided their own draft text to the LLM, which in turn suggested improvements such as corrections of grammatical errors, clearer phrasing, and removal of non-academic expressions. LLMs were also used to inspire possible titles for the paper. While the system provided suggestions, the final title was decided and refined by the authors and is not directly taken from any single LLM output. In addition, LLMs were used as coding assistants during the implementation phase. They provided code completion and debugging suggestions, but all final implementations, experimental design, and validation were carried out and verified by the authors. Importantly, LLMs were NOT used for generating research ideas, designing experiments, or searching and reviewing related work. All conceptual contributions and experimental designs were fully conceived and executed by the authors.

### v3

@@ 5 Math Magic
Hypothesis.
Consider the following recreational arithmetic game, a variant of the Kaprekar routine Kaprekar (1955), which relies on digit manipulation in base 10. Think of a three-digit number in which the hundreds and units digits differ (e.g., abc). Reverse the digits to form a new number (e.g., abc $\rightarrow$ cba), and subtract the smaller of the two numbers from the larger. Then, reverse the result—if it is a two-digit number, prepend a zero (e.g., 67 $\rightarrow$ 067 $\rightarrow$ 760). Add this reversed number to the previous result. Unsurprisingly, such computation always leads to 1089. This deterministic convergence is an example of “mathematical magic” or “math mentalism”—procedures that appear mysterious but are fully explained by the arithmetic structure of decimal digits. The invariance arises from the fact that the subtraction step yields a multiple of 99, and the subsequent reversal–addition step collapses all cases to the constant 1089.
From a cognitive perspective, performing such routines requires humans to encode digits in working memory, apply digit-level transformations (reversing, subtracting, padding), and track intermediate results internally—similar to remembering a poker card during a trick. This process relies on working memory, a lack of which would lead to failure to reproduce these deterministic outcomes when asked to simulate the trick.
@@
Our preliminary experiments show that LLMs can recognize and accurately predict the number 1089, suggesting that this well-known game is likely included in the training data, invalidating this game as an evaluation protocol. To more effectively assess LLMs’ capability for multi-step mental manipulation, we select a more complicated routine based on the Josephus Problem Schumer (2002). In this task, participants are asked to imagine four numbers and perform a sequence of operations, including duplication, rotation, and removal. The full procedure is illustrated in Table 5. Ultimately, only two numbers remain, and mathematical constraints guarantee they are identical. In our experiment, we prompt LLMs to privately select four numbers and mentally execute the sequence of operations. We report the proportion of 150 trials in which the model correctly produced two identical numbers.
Results.
Table 6a reports the accuracy of prompting models to output the two numbers directly. Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family. This finding aligns with results from the number guessing game shown in Table 1, where LLaMA models generate more realistic distributions than other models. Taken together, these findings point to a consistent trend: while some models perform marginally better, current LLMs generally fail to maintain the internal state required for this kind of sequential manipulation.
@@ Results.
We further examine whether CoT prompting improves performance on this task. Table 6b presents the results of prompting models to reason step by step, as well as the performance of o1-like long reasoning models (LRMs). Base models prompted to reason step-by-step achieve 10–30% accuracy—substantially higher than without CoT. DeepSeek-R1 attains 100% accuracy, and other LRMs also perform well. Notably, models also exhibit a strong preference for the number seven, consistent with our number-guessing experiment. For example, 66.7% of o1-Mini’s correct predictions, 46.9% of o3-Mini’s, and 68.5% of o4-Mini’s involve the number seven. Notably, o3-Mini—being least likely to guess 7—achieves a higher accuracy than other two o-series models. These findings suggest that CoT and LRMs can improve accuracy by externalizing intermediate steps, but the success depends on explicit reasoning tokens rather than latent persistence for working memory. The persistence of number preference bias and failure on these tasks suggests that current LLMs struggle with tasks that require sustained internal state and mental manipulation.
6 Discussion
Summary.
@@ 6 Discussion
In this study, we present three carefully designed experiments to investigate whether LLMs exhibit human-like working memory. Across all experiments, the results reveal a consistent pattern: LLMs do not exhibit behavior indicative of a functional working memory. They fail to internally represent or manipulate transient information across multiple reasoning steps, relying instead on the immediate prompt context. Even advanced prompting strategies, such as CoT prompting, yield only marginal improvements on tasks requiring internal state management.
Implications.
The absence of working memory manifests in three ways: unrealistic responses, self-contradictions, and inability to perform mental manipulations. This deficit directly constrains LLM performance on real-world tasks that require internal state maintenance for execution, including real-world planning tasks such as travel planning Xie et al. (2024a); Wang et al. (2025a), scientific inquiry Nathani et al. (2025), and application navigation Xie et al. (2024b); He et al. (2024); Lyu et al. (2025). The challenges are further magnified in multi-agent settings: without working memory, LLM agents quickly lose track in extended dialogues Laban et al. (2025), abandon their initial goals (goal drift Arike et al. (2025)), or mistakenly adopt others’ perspectives as their own (identity drift Choi et al. (2024)). Moreover, for LLM-based multi-agent social simulation, the lack of working memory departs LLMs from real-world human subjects, potentially invalidating the simulation as the behavior is fundamentally different Zhou et al. (2025). In short, the lack of working memory is not just a theoretical concern: it directly undermines reliability, coherence, and validity in applied AI systems. In human cognition, both are necessary: we reason aloud and also rely on a silent working memory buffer to hold commitments, track goals, and compare states. The absence of this buffer in LLMs may explain why they excel at visible reasoning (e.g., think step by step) yet collapse when asked to “think silently.”
Future work.
A natural next step is to explore mechanisms that could grant LLMs intrinsic working memory. While engineering approaches such as external text- or vector-based memories can compensate for some deficits, they do not address the core limitation: LLMs’ inability to sustain internal, latent state over time. We argue that solutions should move beyond external augmentation toward intrinsic mechanisms—architectural innovations, recurrent depth, or hybrid symbolic–neural components—to provide robust working memory. Interpretability studies have shown that specialized attention heads Wang et al. (2023); Olsson et al. (2022) or expert subnetworks Cai et al. (2025) encode distinct functions, hinting at potential internal substrates for working memory. Such development could bridge the gap between superficial token recall and genuine state maintenance, enabling more human-like reasoning, advancing both reliability and cognitive plausibility.